OpenStudy (anonymous):

Help with algebra ? What number must you add to complete the square? x2 + 10x = 29

OpenStudy (anonymous):

2 is an exponent

OpenStudy (retireed):

I never was any good with these. I always forget the process.

OpenStudy (anonymous):

haha same i have like no idea what to do

OpenStudy (retireed):

I think you need to add a half of a * b where a and b are ax^2 + bx so add ab/2 Let me work on it some okay?

OpenStudy (anonymous):

no problem!

OpenStudy (retireed):

No that's not right. I think now you add (ab/2)^2 Trying again....

OpenStudy (anonymous):

the options are 25, 5, 20, 10

OpenStudy (retireed):

Oh that helps!

OpenStudy (retireed):

So x^2 + 10x + ??? = 29 (ab/2)^2 is 1*10/2 squared equals 5^2 or 25 the answer x^2 + 10x + 25 = 29 +25 (x+5) (x+5) = 54 (x+5)^2 = 54 That's what I think it is after I added 25 in order to complete the square.

OpenStudy (anonymous):

Thank you! @retirEEd